Abbot and Costello are transporting a 0.125m3 cubic block along a ramp inclined at 20°. They tied the

block under the midpoint of a 3m, 2kg, pole-arm. Abbot is at the lower end of the pole while Costello

leads the pair up, holding the other end of the pole. Calculate the upward forces exerted by Abbot

and Costello during the transport if the block weighs 1000N.


1
Expert's answer
2021-03-31T07:15:46-0400

Solution.

"V_c=0.125m^3;"

"\alpha=20^o;"

"L=3m;"

"m_r=2kg;"

"w_c=1000N;"

"F-?;"

As a rule of moments we have:

"(w_c+m_rg)cos\alpha\sdot\dfrac{L}{2}=F\sdot L
\implies"

"F=\dfrac{(w_c+m_rg)cos\alpha}{2};"

"F=\dfrac{(1000N+20N)cos20^o}{2}=479N" ;

Aswer: "F=479N."